# Leveraging Prototypical Representations for Mitigating Social Bias without Demographic Information

### v1

@@ Appendix E Full Results
Appendix F Effect of Number of Prototypical Text Pairs
To investigate the effect of the number of pairs, K, on model performance, we conducted experiments with varying K values (1, 2, 4, 8). The results, presented in Table 8, reveal that all K values contributes to the reduction of the TPR-GAP without affecting accuracy. While larger values of K result in more substantial reductions, the incremental improvements become less significant for $K>2$. These findings suggest that a small K may be sufficient for the algorithm.
Occupation Prediction Sentiment Analysis
K
Accuracy $\uparrow$ TPR-GAP $\downarrow$ Accuracy $\uparrow$ TPR-GAP $\downarrow$
Original
$83.43\pm{\scriptstyle 0.08}$ $14.66\pm{\scriptstyle 0.51}$ $79.18\pm{\scriptstyle 0.22}$ $25.34\pm{\scriptstyle 0.82}$
1
$82.85\pm{\scriptstyle 0.18}$ $13.11\pm{\scriptstyle 0.62}$ $77.52\pm{\scriptstyle 1.23}$ $22.69\pm{\scriptstyle 1.81}$
2
$82.55\pm{\scriptstyle 0.23}$ $12.84\pm{\scriptstyle 0.59}$ $77.16\pm{\scriptstyle 1.12}$ $22.21\pm{\scriptstyle 1.08}$
4
$82.32\pm{\scriptstyle 0.13}$ $12.29\pm{\scriptstyle 0.32}$ $77.20\pm{\scriptstyle 1.17}$ $21.72\pm{\scriptstyle 0.82}$
8
$82.29\pm{\scriptstyle 0.30}$ $12.20\pm{\scriptstyle 0.25}$ $77.47\pm{\scriptstyle 0.94}$ $22.17\pm{\scriptstyle 1.17}$

### v2

@@ Appendix C Full Results
C.1 Accuracy Figures
(a) Occupation Prediction
(b) Sentiment Analysis
Figure 3: Effect of bias mitigation methods on Accuracy with varying labeled data sizes.
Male Prototypical Texts Female Prototypical Texts
This is a biography about a male. This is a biography about a female.
A man who excelled in his field. A woman who excelled in her field.
He is known for his achievements in various industries. She is known for her achievements in various industries.
A prominent male figure in history. A prominent female figure in history.
His career and accomplishments are well-regarded. Her career and accomplishments are well-regarded.
This biography focuses on the life of a distinguished man. This biography focuses on the life of a distinguished woman.
An influential male individual. An influential female individual.
He made significant contributions to his profession. She made significant contributions to her profession.
This is a story about a man who shaped his industry. This is a story about a woman who shaped her industry.
His impact on his field is noteworthy. Her impact on her field is noteworthy.
AAE Prototypical Texts SAE Prototypical Texts
This tweet reflects a [sentiment] from a white writer. This tweet reflects a [sentiment] sentiment from a black writer.
A tweet expressing a [sentiment] moment by a white individual. A tweet expressing a [sentiment] moment by a black individual.
A [sentiment] viewpoint shared by a writer using Standart American English. A [sentiment] viewpoint shared by a writer using African American English.
This post, written in standard English, conveys [sentiment] from a white perspective. This post, written in AAE, conveys [sentiment] from a black perspective.
A message filled with [sentiment] from a white communicator. A message filled with [sentiment] from a black communicator.
A white person shares their [sentiment] thoughts in this tweet. A black person shares their [sentiment] thoughts in this tweet.
This is an example of a tweet with [sentiment] in white sociolect. This is an example of a tweet with [sentiment] sentiment in AAE.
A tweet written by a white speaker that conveys [sentiment]. A tweet written by a black speaker that conveys [sentiment].
This post by a white individual radiates [sentiment] and [sentiment]. This post by a black individual radiates [sentiment] and [sentiment].
A [sentiment] perspective presented by a writer using white sociolect. A [sentiment] perspective presented by a writer using African American English.
Table 8: Pre-defined Representations for AAE and SAE Tweet Texts
Figure 4: The prompt for generating prototypical text pairs.